OpenStudy (anonymous):

Two vectors A and B have precisely equal magnitudes. In order for the magnitude of A+ B to be 110 times larger than the magnitude of A-B , what must be the angle between them?

OpenStudy (anonymous):

Thanks for pointing that out :)

OpenStudy (amistre64):

yeah, it helps when trying to solve the problem to know what to solve :)

OpenStudy (amistre64):

Are these vectors in any particular dimension? R^n ?

OpenStudy (anonymous):

no particular direction but assume its positive.

OpenStudy (amistre64):

\[A=<a_1,a_2,a_3,...,a_n>;\ |A|=\sqrt{(a_1)^2+(a_2)^2+(a_3)^2+...+(a_n)^2} \] \[B=<b_1,b_2,b_3,...,b_n>;\ |B|=\sqrt{(b_1)^2+(b_2)^2+(b_3)^2+...+(b_n)^2}\] so far

OpenStudy (amistre64):

\[A+B=<a_1+b_1,a_2+b_2,a_3+b_3,...,a_n+b_n>\] \[ |A+B|=\sqrt{(a_1+b_1)^2+(a_2+b_2)^2+(a_3+b_3)^2+...+(a_n+b_n)^2}\] \[A-B=<a_1-b_1,a_2-b_2,a_3-b_3,...,a_n-b_n>\] \[ |A+B|=\sqrt{(a_1-b_1)^2+(a_2-b_2)^2+(a_3-b_3)^2+...+(a_n-b_n)^2}\] this might get messy

OpenStudy (amistre64):

especially if i forget how to type ... that last ones |A-B|

OpenStudy (amistre64):

but this is going towards angle between them; so i really should focus on that ...

OpenStudy (amistre64):

\[A.B=a_1b_1+a_2b_2+a_3b_3+...+a_nb_n\] and we determined the magnitudes already soo; the formula for the angle between them is: \[cos(t)=\frac{A.B}{|A||B|}\] \[t=cos^{-1}\frac{A.B}{|A||B|}\] \[t=cos^{-1}\frac{a_1b_1+a_2b_2+a_3b_3+...+a_nb_n}{110((a_1-b_1)^2+(a_2-b_2)^2+(a_3-b_3)^2+...+(a_n-b_n)^2)}\]

OpenStudy (amistre64):

you sure this thing is any dimensional?

OpenStudy (anonymous):

this is motion in 2 dimensions, if this answers your question.

OpenStudy (amistre64):

well, it does; makes that contraption up there a bit smaller :)

OpenStudy (anonymous):

i think so too.

OpenStudy (amistre64):

\[t=cos^{-1}\frac{a_1b_1+a_2b_2}{110((a_1-b_1)^2+(a_2-b_2)^2)}\] \[110t=cos^{-1}\frac{a_1b_1+a_2b_2}{(a_1)^2-2a_1b_1+(b_1)^2+(a_2)^2-2a_2b_2+(b_2)^2}\] how that helps tho I cant tell; might have to wolf it to see how or if it simplifies

OpenStudy (anonymous):

thanks, i'll solve and let you know.

OpenStudy (amistre64):

 this might be an easier thing to construct and see a solution with

OpenStudy (amistre64):

|A+B| = 110|A-B|